### v2

# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

Abstract.
In motion generation, controllability as well as generation quality and speed is becoming more and more important. There are various motion editing tasks, such as in-betweening, upper body editing, and path-following, but existing methods perform motion editing with a data-space diffusion model, which is slow in inference compared to a latent diffusion model. In this paper, we propose MoLA, which provides fast and high-quality motion generation and also can deal with multiple editing tasks in a single framework. For high-quality and fast generation, we employ a variational autoencoder and latent diffusion model, and improve the performance with adversarial training. In addition, we apply a training-free guided generation framework to achieve various editing tasks with motion control inputs. We quantitatively show the effectiveness of adversarial learning in text-to-motion generation, and demonstrate the applicability of our editing framework to multiple editing tasks in the motion domain. We will make our code publicly available.
Motion generation, Latent diffusion model, Adversarial training, Guided generation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computer visionccs: Computing methodologies Motion processing
Figure 1. MoLA achieves fast and high-quality human motion generation given textual descriptions while enabling motion editing applications. With MoLA, we can deal with various types of motion editing tasks in a single framework. The above images show the standard text-to-motion generation (as a no-control input case) and three motion editing tasks (with control inputs): the control input for the start-end positions and motion in-between task (left), for the lower body positions and upper body editing task (center), for the motion trajectory and path-following generation task (right).
1. Introduction
Fast generation
High-quality generation
Training-free editing
MDM-based methods (Tevet et al., 2023; Karunratanakul et al., 2023b; Xie et al., 2023; Karunratanakul et al., 2023a) - -
Motion Latent Diffusion (MLD) (Chen et al., 2023) - -
VQ-based methods (Zhang et al., 2023b; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024)   -
MoLA (ours)
Table 1. Comparison of three properties satisfied by MoLA and other text-to-motion methods. $\dagger$ Some editing tasks can be done in a training-free manner, but there are editing tasks that cannot be done.
Motion generation is a fundamental task in multimedia production and computer animation, and it can be challenging due to the diversity of possible motions and the difficulty and cost of obtaining high-quality data. Generating human motions from textual descriptions, i.e., text-to-motion generation, is an emerging task that may play an important role in a broad range of applications such as games, film, the metaverse, and virtual reality. Over the past few years, text-to-motion has generated intensive research interest (Tevet et al., 2023; Wang et al., 2023; Zhong et al., 2023; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024; Zhang et al., 2023b; Chen et al., 2023; Zhang et al., 2024). The existing methods of text-to-motion encompass several different approaches. For example, some utilize vector quantization and have achieved high generation performance (Zhang et al., 2023b; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024). Inspired by the development of diffusion models and their convenience, several diffusion-based text-to-motion methods have also been developed to not only generate motion but also to perform editing of the motion samples (Tevet et al., 2023; Zhang et al., 2024). In addition, Chen et al. adopted a latent diffusion model (Rombach et al., 2022) to enable fast motion generation (Chen et al., 2023). Alternatively, other methods for motion editing tasks utilize a pre-trained text-to-motion model as a generative prior, which has attached attention due to its applicability (Song et al., 2023b; Karunratanakul et al., 2023b; Xie et al., 2023; Karunratanakul et al., 2023a).
Ideally, (text-to-)motion generation models should be able to 1) perform fast generation, 2) achieve a high generation performance, and 3) perform editing motions. However, it is difficult for the existing approaches to meet all three requirements. The VQ-based text-to-motion methods (Zhang et al., 2023b; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024) require additional training for each editing task, as discussed in (Pinyoanuntapong et al., 2024). In other words, the VQ-based methods cannot control arbitrary joints in a training-free manner. Requiring extra training per editing task harms the applicability of motion generation models to production. Existing motion editing methods (Song et al., 2023b; Karunratanakul et al., 2023b; Xie et al., 2023; Karunratanakul et al., 2023a) utilize diffusion-based methods in the data-space, such as Motion Diffusion Model (MDM) (Tevet et al., 2023), as a generative prior, but the generation speed of diffusion in the data-space is not fast. In contrast, the existing latent diffusion-based motion generation method (Chen et al., 2023) performs fast generation, but motion editing based on continuous latent diffusion has not been explored yet.
In this paper, we aim to develop a text-to-motion model that achieves 1) fast generation, 2) a high generation performance, and 3) multiple motion editing in a training-free manner. We summarize the target properties satisfied by our model in Table 1 in comparison with the existing text-to-motion studies. To achieve this goal, we employ a motion variational autoencoder (VAE) to learn a low-dimensional latent representation for diverse human motion sequences and a latent diffusion model using text condition as in (Chen et al., 2023). Here, we combine a motion VAE with adversarial training to achieve better reconstruction/generation. It has been verified that adversarial training improves the performance of VAE in the image and audio domains (Esser et al., 2021; Iashin and Rahtu, 2021; Rombach et al., 2022; Liu et al., 2023), but not yet in motion generation. To the best of our knowledge, our model is the first latent-space motion generation model featuring a VAE is trained with adversarial training. In addition, we introduce training-free guided diffusion (He et al., 2024) into diffusion sampling on latent space to enable multiple motion editing tasks. We empirically show that adversarial training increases the quality on two widely utilized datasets. We also demonstrate that training-free guided diffusion can enable multiple motion editing tasks such as in-betweening, upper body editing, and path-following.
The main contributions of this paper are as follows:
  • We propose a new text-to-motion model that can perform fast and high-quality generation and can deal with various types of motion editing tasks (motion in-betweening, upper body editing, and path following) in a training-free manner. To the best of our knowledge, this is the first study to demonstrate motion editing with latent diffusion.
  • We quantitatively show that introducing adversarial training to a motion VAE improves reconstruction and generation performance on text-to-motion. To the best of our knowledge, this is the first latent-space motion generation model that introduces adversarial training.
2. Related Work
2.1. Deep Generative Modeling
The progress in generative modeling has developed rapidly over the past decade (Kingma and Welling, 2013; Goodfellow et al., 2014; Sohl-Dickstein et al., 2015). In recent years, diffusion models have been actively explored, particularly in the image domain. They have achieved state-of-the-art performance not only in image generation (Gao et al., 2023; Peebles and Xie, 2023) but also in advanced generative tasks such as text-to-image (Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022), image inpainting (Lugmayr et al., 2022; Wang et al., 2022), and guided generation (Yu et al., 2023; Zhang et al., 2023a). A key approach for reducing inference costs is to reduce the essential dimensionality of the target data by learning its compressed representation (Rombach et al., 2022; Gao et al., 2023). This approach involves two stages: learning a VAE to acquire latent representations and then training a diffusion model on the low-dimensional representations.
A pioneering work that incorporates such a two-stage training scheme into generative modeling is VQ-VAE, which aims to compress the target data in a discrete manner (van den Oord et al., 2017). Several subsequent works have followed and expanded this discrete approach (Razavi et al., 2019; Esser et al., 2021; Lee et al., 2022). Esser et al. attempted to improve the reconstruction fidelity of VQ-VAE through adversarial training (Esser et al., 2021), which has been adopted by many subsequent works (Lee et al., 2022; Gu et al., 2022; Chang et al., 2022, 2023). As another advancement in the stage 1 model, Lee et al. generalized the VQ operation to residual vector quantization (RVQ), precisely approximating the encoded feature as a sequence of code stacks (Lee et al., 2022). In contrast, VQ-Diffusion adopts a discrete diffusion model that allows the input of a text prompt for learning the discrete latent, thereby achieving text-to-image generation (Gu et al., 2022). Mask-prediction training schemes with confidence-based iterative decoding have been proposed for the stage 2 (Chang et al., 2022, 2023).
2.2. Training-free Guided Generation
Starting with works such as (Song et al., 2021b; Meng et al., 2022), there have been many attempts to utilize pre-trained diffusion models as rich priors for various conditional generation tasks without additional training (i.e., training-free(Chung et al., 2023; Wang et al., 2022; Song et al., 2023b, a). The fundamental idea behind these approaches is to guide the intermediate samples of the diffusion model during the generation so that the final generated output satisfies the given conditions.
With increasing attention on the practical usefulness of latent diffusion models (LDMs), some recent works have focused on methods capable of dealing with LDMs. FreeDoM (Yu et al., 2023) performs a variety of conditional generation tasks in a unified framework by designing a task-specific loss function and utilizing the loss during guided sampling. Approaches such as PSLD (Rout et al., 2023) and MPGD (He et al., 2024) are attempting to improve on this even more, achieving higher performance and faster guided generation.
2.3. Motion Generation
Text-to-motion generation technology has made rapid progress with the development of neural networks (Ahuja and Morency, 2019; Petrovich et al., 2022; Guo et al., 2022; Tevet et al., 2022; Petrovich et al., 2023). In particular, diffusion models and VQ-based models have significantly advanced the field of motion generation.
MDM (Tevet et al., 2023) and MotionDiffuse (Zhang et al., 2024) adopt diffusion models for motion generation, which leads to better performance in terms of generation quality. However, these methods directly apply diffusion processes to raw motion sequences, thus resulting in slow generation. MLD (Chen et al., 2023) mitigates this issue by adopting a diffusion model in a low-dimensional latent space provided by a VAE trained on motion data, inspired by latent diffusion models (Rombach et al., 2022).
VQ-based models have been studied as well, inspired by the success of VQ-based models in image generation (Gu et al., 2022; Yu et al., 2022; Chang et al., 2023). In this approach, a VQ-VAE model is first trained on motion data to acquire discrete motion representations, and deep generative models are then applied to generate sequences of discrete representations (also called tokens). T2M-GPT (Zhang et al., 2023b), AttT2M (Zhong et al., 2023), and MotionGPT (Jiang et al., 2023) utilize autoregressive (AR) models to generate motion tokens. However, AR models are slow in inference because motion tokens are generated sequentially. To address this issue, M2DM (Kong et al., 2023) and DiverseMotion (Lou et al., 2023) apply discrete diffusion models to motion tokens in a latent space, whereas MMM (Pinyoanuntapong et al., 2024) adopts a mask prediction model. In another line of research, MoMask (Guo et al., 2023) leverages the RVQ technique to reduce errors caused by the VQ process.
In this paper, we adopt a continuous diffusion model to perform multiple editing tasks with a single framework (Section 3.4). To achieve high-fidelity and fast generation, we boost the performance of a VAE model with adversarial training and apply a diffusion model in the latent space provided by the trained VAE.
2.4. Motion Editing
Motion editing has attracted much research interest as well. MDM (Tevet et al., 2023) demonstrated upper body editing and motion in-betweening by applying diffusion inpainting to motion data in both the spatial and temporal domains. LGD (Song et al., 2023b) demonstrated path-following motion generation with a guided diffusion that utilizes multiple samples from a suitable distribution to reduce bias. GMD (Karunratanakul et al., 2023b) guides the position of the root joint to control motion trajectories. OmniControl (Xie et al., 2023) controls any joints at any time by guiding a pre-trained motion diffusion model with an analytic function. DNO (Karunratanakul et al., 2023a) optimizes the diffusion latent noise of a pre-trained text-to-motion model with user-provided criteria in the motion space and achieves multiple editing tasks. However, these methods employ data-space diffusion models similar to MDM and have not been demonstrated with a latent diffusion model. Recently, MMM (Pinyoanuntapong et al., 2024) demonstrated motion editing by placing masked tokens in the place that needs editing and applying the mask prediction framework. However, it cannot control an arbitrary body part without additional model training, and the authors trained separate upper-body and lower-body generative models to perform upper-body editing.
In this paper, we employ a latent diffusion model to achieve fast generation and apply a training-free guided generation framework that can deal with multiple editing tasks without additional model training.
3. Method
3.1. Overall Framework of MoLA
The goal of this study is to develop a framework for fast and high-quality text-guided motion generation and to deal with multiple control tasks in a training-free way. To achieve this, we propose the following training and inference methods:
  • Stage 1 training: We first train a motion VAE enhanced by adversarial training to achieve high reconstruction/generation performance. We further propose replacing the decoder for adversarial training with a SAN-based one to boost our model performance (Section 3.2).
  • Stage 2 training: To reduce computational complexity while simultaneously enabling high-quality text-driven motion generation, we train a text-conditioned diffusion model on the low-dimensional latent space learned by stage 1, i.e., a latent diffusion model (Section 3.3).
  • Guided generation: Adopting training-free guided generation in inference enables multiple editing functions required in motion generation without additional training (Section 3.4).
The outline of our text-to-motion generation model, MoLA, is shown in Figure 2.
Figure 2. Overall architecture of MoLA. Stage 1: A motion VAE enhanced by adversarial training learns a low-dimensional latent representation for diverse human motion sequences. Stage 2: A text-conditioned latent diffusion model reduces computational costs and enables high-quality text-to-motion generation by utilizing the latent representation obtained in stage 1. Guided generation: During the inference of the stage 2 model, a training-free guided generation method minimizes a loss function $L$, which can be flexibly set depending on the task, and multiple motion editing tasks can also be performed in a single framework.
3.2. Stage 1: Continuous Motion Latent Representation with Adversarial Training
3.2.1. Learning a rich motion latent representation with VAE-GAN
We propose a motion variational autoencoder (VAE), enhanced by adversarial training, to learn a low-dimensional latent representation for diverse human motion sequences. Assume an observed motion $\bm{x}\in\mathbb{R}^{N\times L}$, where $N$ and $L$ denote raw motion data dimension per frame and motion length, respectively. To learn such a latent representation, we first define a $d_{z}$-dimensional latent variable $\bm{z}\in\mathbb{R}^{d_{z}}$, which is assumed to generate data sample $\bm{x}$. The generative process is modeled as $\bm{x}\sim p_{\bm{\psi}}(\bm{x}|\bm{z})$ with a prior $p(\bm{z})$. The prior is assumed to be a standard Gaussian distribution, i.e., $p(\bm{z})=\mathcal{N}(\bm{0},\bm{I})$. We model the conditional distribution as a Gaussian distribution: $p_{\bm{\psi}}(\bm{x}|\bm{z})=\mathcal{N}(g_{\bm{\psi}}(\bm{z}),\sigma^{2}\bm{I})$ with $g_{\bm{\psi}}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{N\times L}$ and $\sigma^{2}\in\mathbb{R}_{+}$, where $\mathbb{R}_{+}$ indicates the set of all positive real numbers. As in a usual VAE, we introduce an approximated posterior, which is modeled by a Gaussian distribution as $q_{\bm{\eta}}(\bm{z}|\bm{x})=\mathcal{N}(\bm{\mu}_{\bm{\eta}}(\bm{x}),\mathrm{diag}(\bm{\sigma}_{\bm{\eta}}(\bm{x})))$, with $\bm{\mu}_{\bm{\eta}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{z}}$ and $\bm{\sigma}_{\bm{\eta}}:\mathbb{R}^{N\times L}\to\mathbb{R}_{+}^{d_{z}}$. As a result, the VAE consists of an encoder and a decoder, parameterized by $\bm{\eta}$ and $\bm{\psi}$, respectively. We construct the autoencoder based on a transformer-based architecture (Petrovich et al., 2021; Chen et al., 2023). The objective function for the VAE is formulated as the negative evidence lower bound (negative ELBO) per sample $\bm{x}$, which is a summation of mean squared error and KL regularization terms:
(1) $\mathcal{J}_{\text{VAE}}(\bm{\psi},\bm{\eta};\bm{x})=$ $~{}\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[\|\bm{x}-g_{\bm{\psi}}(\bm{z})\|_{2}^{2}]+\lambda_{\text{reg}}D_{KL}(q_{\bm{\eta}}(\bm{z}|\bm{x})\|p(\bm{z})),$
where $\lambda_{\text{reg}}$ is a reciprocal of $\sigma^{2}$ (treated as a hyperparameter in this work). The encoder $q_{\bm{\eta}}$ can be trained to produce a low-dimensional latent representation, and the decoder $g_{\bm{\psi}}$ can also be trained to accurately reconstruct the input motion sequences from the latent representations.
To achieve high-quality generation, we need to push the limits of compression. Hence, we propose incorporating adversarial training into the motion VAE. More specifically, we introduce a discriminator, denoted as $f_{\bm{\phi}}:\mathbb{R}^{N\times L}\to\mathbb{R}$, that aims to distinguish real and reconstructed motions. The adversarial training is formulated as a two-player optimization between the VAE and the discriminator. The discriminator is trained by the maximization of $\mathbb{E}_{p(\bm{x})}\mathcal{L}_{\text{GAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})$ with respect to $\bm{\phi}$, where
$\mathcal{L}_{\text{GAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})=$ $~{}\min\{0,-1+f_{\bm{\phi}}(\bm{x})\}$
(2) $+\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}\left[\min\{0,-1-f_{\bm{\phi}}(g_{\bm{\psi}}(\bm{z}))\}\right].$
We formulate the overall loss for the VAE as a combined objective of the negative ELBO and adversarial loss, as
(3) $\min_{\phi,\psi}\mathbb{E}_{p(x)}\left[\mathcal{J}_{\text{VAE}}(\bm{\psi},\bm{\eta};\bm{x})+\lambda_{\text{adv}}\mathcal{J}_{\text{GAN}}(\bm{\psi},\bm{\eta};\bm{\phi},\bm{x})\right],$
where $\lambda_{\text{adv}}$ is a positive scalar that adjusts the balance between the two terms, and
(4) $\mathcal{J}_{\text{GAN}}(\bm{\psi},\bm{\eta};\bm{\phi},\bm{x})=$ $~{}-\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[f_{\bm{\phi}}(g_{\bm{\psi}}(\bm{z}))].$
We train both the VAE and the discriminator with Equations (2) and (3) in an alternating way.
3.2.2. From GAN- to SAN-based discriminator
We apply the SAN framework (Takida et al., 2024) to further enhance the motion VAE based on a prior report showing that a model with a SAN-based discriminator has achieved SOTA performance in image generation. Takida et al. (Takida et al., 2024) incorporated a sliced optimal transport perspective into a general GAN. Following this work, we first decompose the discriminator $f_{\bm{\phi}}$ into the last linear layer and the remaining neural part, denoted as $\bm{w}\in\mathbb{R}^{d_{w}}$ and $h_{\bm{\varphi}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{w}}$ with $d_{w}\in\mathbb{N}$, respectively. This decomposition provides an interpretation of the discriminator: the neural function $h_{\bm{\varphi}}$ maps motion sequences to non-linear features, and then the linear layer $\bm{w}$ projects them into scalars. As a preliminary step for applying SAN to the adversarial training in Section 3.2.1, we normalize $w$ using its norm, resulting in $\bm{\omega}=\bm{w}/\|\bm{w}\|_{2}$, which indicates the direction of the projecting. Now, the discriminator is represented in the form of an inner product as $f_{\bm{\phi}}(\bm{x})=\bm{\omega}^{\top}h_{\varphi}(\bm{x})$, and its parameter is $\bm{\phi}=\{\bm{\varphi},\bm{\omega}\}$.
The prior work has shown that the discriminator obtained from the optimal solution of $\bm{\omega}$ in the hinge loss (2) does not guarantee gradients that make the generated distribution close to the data distribution. To address this issue, we adopt the SAN maximization problem instead of Equation (2). Specifically, we optimize the neural part $\bm{\varphi}$ with the original hinge loss, while applying the Wasserstein GAN loss (Arjovsky et al., 2017) to the direction $\bm{\omega}$. The modified maximization objective is formulated as
$\mathcal{L}_{\text{SAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})$ $=\mathcal{L}_{\text{GAN}}(\{\bm{\varphi},\bm{\omega}^{-}\};\bm{\psi},\bm{\eta},\bm{x})$
(5) $+\bm{\omega}^{\top}\left(h_{\bm{\varphi}^{-}}(\bm{x})-\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[h_{\bm{\varphi}^{-}}(g_{\bm{\psi}}(\bm{z}))]\right),$
where $(\cdot)^{-}$ indicates a stop gradient operator. The second term in Equation (5) induces the direction that best discriminates between the real and generated sample sets in the feature space. We employ the same minimization objective as defined in Equation (3).
3.3. Stage 2: Motion Latent Diffusion
In this section, we train a text-conditioned diffusion model on the low-dimensional motion latent space obtained by the autoencoder learned in the stage 1 (Section 3.2). Using the trained model, we perform motion generation conditioned on text. First, we define a time-dependent sequence $\bm{z}=\bm{z}_{0},\bm{z}_{1},\dots,\bm{z}_{t},\dots,\bm{z}_{T}\in\mathbb{R}^{d_{z}}$ starting from the VAE encoder output $\bm{z}\sim q_{\bm{\eta}}(\bm{z}|\bm{x})$, which is derived from the following Markov diffusion process in the latent space:
(6) $q(\bm{z}_{t}|\bm{z}_{t-1})=\mathcal{N}(\sqrt{\alpha}_{t}\bm{z}_{t-1},(1-\alpha_{t})\bm{I}),$
where $T$ 0 and the constant $\alpha_{t}\in(0,1)$ is a pre-defined noise-scheduling parameter that determines the forward process. The forward process allows for the sampling of $\bm{z}_{t}$ at an arbitrary time step $t$ in a closed form: $\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\bar{\alpha}_{t}\coloneq\prod_{s=1}^{t}\alpha_{s}$ and $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$.
The diffusion model is trained to reverse the diffusion process described in Equation (6), i.e., to perform denoising. Specifically, the model learns to estimate the noise component $\epsilon$ in the noisy sample $\bm{z}_{t}$ as in the following objective:
(7) $\mathcal{J}_{\text{LDM}}(\theta)=\mathbb{E}_{\bm{z}_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},t)\|_{2}^{2}\right],$
where $\theta$ is the model parameter to be trained, and the time step $t$ is uniformly sampled from $\{1,\dots,T\}$.
As our goal is text-to-motion generation, our interest is in the conditional distribution $p(\bm{z}|\bm{c})$ conditioned on the text prompt $\bm{c}$. Here, similar to many text-conditioned latent diffusion models (Rombach et al., 2022; Chen et al., 2023), we train the conditional model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ conditioned on the output of a text encoder $\tau(\bm{c})$, using the following objective function:
(8) $\mathcal{J}_{\text{cLDM}}(\theta)=\mathbb{E}_{\{\bm{z}_{0},\bm{c}\},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))\|_{2}^{2}\right],$
where $\bm{z}_{0}$ and $\bm{c}$ are drawn from the joint empirical distribution. In addition, following many other works, we adopt classifier-free guidance (Ho and Salimans, 2022) and train the model unconditionally i.e., without a text prompt, with a certain probability during training.
During the inference (text-to-motion generation), the trained diffusion model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ is employed to generate $\bm{z}_{0}$ through a denoising process conditioned on the text prompt $\bm{c}$. We adopt the sampling scheme of DDIM (Song et al., 2021a), in which each sampling step is defined as:
$\bm{z}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\bigg{(}\frac{\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))}{\sqrt{\bar{\alpha}_{t}}}\bigg{)}$
(9) $\ +\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))+\sigma_{t}\epsilon,$
where $\sigma_{t}>0$ determines the stochasticity of the sampling process, and the sampling process becomes deterministic when $\sigma_{t}=0$. The part $(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c})))/\sqrt{\bar{\alpha}_{t}}$ in the first term corresponds to a direct estimate of the clean latent $\bm{z}_{0}$ from the noisy sample $\bm{z}_{t}$ using the diffusion model based on Tweedie’s formula (Efron, 2011); this estimate is denoted as $\bm{z}_{0|t}$.
3.4. Controllable Motion Generation on Latent Diffusion Sampling
In this section, we present a guided generation framework that leverages the pre-trained motion latent diffusion model (Section 3.3) for conditional motion generation and editing tasks without extra training. Major training-free methods such as DPS (Chung et al., 2023), FreeDoM (Yu et al., 2023), and MPGD (He et al., 2024) are based on the fact that the conditional score function can be decomposed into two additive terms: the unconditional score function and the log-likelihood term. Specifically, for a new condition $\bm{y}$, we have
(10) $\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t}|\bm{c},\bm{y})=\nabla_{\bm{z}_{t}}\log p(\bm{z}_{t}|\bm{c})+\nabla_{\bm{z}_{t}}\log p(\bm{y}|\bm{z}_{t},\bm{c}),$
as derived from Bayes’ rule. The conditional generation based on the aforementioned property can be regarded as a sequential procedure implemented as follows. First, from $\bm{z}_{t}$, a denoised sample $\bm{z}_{t-1}$ is obtained by the sampling step in Equation (9), without considering the given condition $\bm{y}$. Subsequently, $\bm{z}_{t-1}$ is further updated using the gradient with respect to $\bm{z}_{t}$ for the log-likelihood term.
The log-likelihood term is defined for noisy sample $\bm{z}_{t}$. In the classifier-guided diffusion (Song et al., 2021b), time-dependent classifiers for $\bm{z}_{t}$ have to be trained, which requires additional training. In contrast, in training-free methods such as DPS, FreeDoM, and MPGD, this term is approximated with the current clean data estimate $\bm{z}_{0|t}$ and a loss function $L(\bm{x};\bm{y})$ defined for clean data. This loss function can be flexibly set depending on the task. For example, in inverse problems of the form $\bm{y}=\mathcal{A}(\bm{x})$, where $\mathcal{A}$ is a differentiable function with respect to $\bm{x}$, the loss function can be set as $\|\bm{y}-\mathcal{A}(\bm{x}_{0|t})\|_{2}^{2}$, where $\bm{x}_{0|t}=g_{\bm{\psi}}(\bm{z}_{0|t})$ is a clean data estimate in the original data domain and obtained through the VAE decoder. In FreeDoM’s FaceID-guided generation, the $\ell_{2}$ distance of features obtained by inputting facial images into a facial recognition model is used as the loss.
Here, we adopt MPGD (He et al., 2024), a fast yet high-quality guidance method applicable to latent diffusion models. Following the denoising step, it updates the denoised sample for the loss function $L(\bm{x};\bm{y})$ as follows:
(11) $\bm{z}_{t-1}\leftarrow\bm{z}_{t-1}-\rho_{t}\sqrt{\bar{\alpha}_{t-1}}\nabla_{\bm{z}_{0|t}}L(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y}),$
where $\rho_{t}$ is a time-dependent step size parameter.
More specifically, in editing motion tasks we generate motions to match given specific poses or a trajectory control signal. To deal with various motion editing tasks in this framework, the loss functions are designed as follows:
(12) $L(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y})=\Sigma_{n}\Sigma_{l}m_{nl}\|R(g_{\bm{\psi}}(\bm{z}_{0|t}))_{nl}-y_{nl}\|_{2},$
where $n$ and $l$ are indices of joint and frame, respectively, and $m_{nl}$ is a binary value indicating whether the control position $y_{nl}$ contains a valid value at frame $l$ for joint $n$, and $R(\cdot)$ is a function to convert the motion features including the joint’s local positions to global absolute locations. We set $L$ for the guided generation to measure the distance between desired constraints $\bm{y}$ and the joint locations of the generated motion. Target locations as constraint $\bm{y}$ can be specified for any subset of joints in any subset of motion frames. [^1] As examples of motion editing, if $\bm{y}$ is given as the start-end positions, we can handle the motion in-betweening. If $\bm{y}$ is given as the lower body positions, we can edit the upper body corresponding to the lower one. If $\bm{y}$ is given as the pelvis trajectory, it corresponds to the path-following task. We leave the task details to Section 4.5. Note that the guided generation framework in Equations (11) and (12) has the potential to generate motion while dealing with a variety of time and spatial constraints not limited to these three task examples. Editing a generated motion to match specific poses or to follow a specific trajectory can be achieved by minimizing $L$ with the update rule in Equation (11).
4. Experiments
4.1. Datasets
We conduct empirical evaluations on two widely used datasets for text-to-motion generation: HumanML3D (Guo et al., 2022) and KIT Motion-Language (KIT-ML) (Plappert et al., 2016). HumanML3D contains 14,616 human motions from the AMASS (Mahmood et al., 2019) and HumanAct12 (Guo et al., 2020) datasets and 44,970 text descriptions. KIT-ML contains 3,911 human motion sequences from the KIT (Mandery et al., 2015) and CMU (CMU, 2003) datasets and 6,278 textual descriptions. Both datasets are split into training, validation, and test sets with proportions of 80%, 5% and 15%, respectively.
Dataset Stage 1 model rFID $\downarrow$
HumanML3D VAE 0.206
VAE + GAN 0.179
VAE + SAN 0.141
KIT-ML VAE 0.311
VAE + GAN 0.293
VAE + SAN 0.280
Table 2. Analysis of motion reconstruction performance (training stage 1 with VAE) on HumanML3D and KIT-ML
4.2. Implementation Details
In our experiment, we train two different types of models, as follows. (i) We train the motion VAE and discriminator $f_{\bm{\psi}}$ described in Section 3.2 and use these stage 1 models to train the stage 2 model $\epsilon_{\theta}$. (ii) We train the motion VQ-VAE in T2M-GPT (Zhang et al., 2023b) and the discriminator, and train the transformer model in the latter stage to generate code indices conditioned on the text description. Case (ii) is intended to verify whether adversarial training is also effective in a VQ-based text-to-motion framework.
In case (i), the motion VAE architecture is based on the VAE part in MLD (Chen et al., 2023). The encoder $q_{\bm{\eta}}$ and decoder $g_{\bm{\psi}}$ of the VAE model all consist of nine layers and four heads with skip connection. For adversarial training, we use the discriminator $f_{\bm{\phi}}$ whose architecture is composed of 1D convolution, residual block, and Leaky ReLU. We set $\lambda_{\text{adv}}=1.0\times 10^{-3}$ as the weight for adversarial training. All our models are trained for 6K epochs with the AdamW optimizer using the batch size of 128 and a cosine-annealing learning rate of $1.0\times 10^{-4}$ with a warm-up. All other settings, including those for the stage 2 model $\epsilon_{\theta}$ training and text embedding $\tau(\bm{c})$, follow the official MLD implementation.
In case (ii), we use motion VQ-VAE whose codebook size is set to $512\times 512$. We utilize the AdamW optimizer, batch size of 256, and exponential moving constant of 0.99. For adversarial training, we use the discriminator $f_{\bm{\phi}}$ whose architecture is composed of 1D convolution, residual block, and Leaky ReLU, the same as in case (i). We train the first 200K iterations with a learning rate of $2.0\times 10^{-4}$, and 100K with a learning rate of $1.0\times 10^{-5}$. All other settings follow the official T2M-GPT implementation.
Dataset Stage 1 model rFID $\downarrow$
HumanML3D VQ-VAE 0.065
VQ-VAE + GAN 0.057
VQ-VAE + SAN 0.051
KIT-ML VQ-VAE 0.665
VQ-VAE + GAN 0.377
VQ-VAE + SAN 0.299
Table 3. Analysis of motion reconstruction performance (training stage 1 with VQ-VAE) on HumanML3D and KIT-ML
Category Method R-Precision $\uparrow$ FID $\downarrow$ MMDist $\downarrow$ Diversity $\rightarrow$ MModality $\uparrow$
Top-1 Top-2 Top-3
N/A Real motion data $0.511^{\pm.003}$ $0.703^{\pm.003}$ $0.797^{\pm.002}$ $0.002^{\pm.000}$ $2.974^{\pm.008}$ $9.503^{\pm.065}$
Discrete M2DM (Kong et al., 2023) $0.497^{\pm.003}$ $0.682^{\pm.002}$ $0.763^{\pm.003}$ $0.352^{\pm.005}$ $3.134^{\pm.010}$ $9.926^{\pm.073}$ $3.587^{\pm.072}$
AttT2M (Zhong et al., 2023) $0.499^{\pm.003}$ $0.690^{\pm.002}$ $0.786^{\pm.002}$ $0.112^{\pm.006}$ $3.038^{\pm.007}$ $9.700^{\pm.090}$ $2.452^{\pm.051}$
T2M-GPT (Zhang et al., 2023b) $0.492^{\pm.003}$ $0.679^{\pm.002}$ $0.775^{\pm.002}$ $0.141^{\pm.005}$ $3.121^{\pm.009}$ $9.722^{\pm.081}$ $1.831^{\pm.048}$
MoMask (Guo et al., 2023) $0.521^{\pm.002}$ $0.713^{\pm.002}$ $0.807^{\pm.002}$ $0.045^{\pm.002}$ $2.958^{\pm.008}$ $1.241^{\pm.040}$
DiverseMotion (Lou et al., 2023) $0.496^{\pm.004}$ $0.687^{\pm.004}$ $0.783^{\pm.003}$ $0.070^{\pm.004}$ $3.063^{\pm.011}$ $9.551^{\pm.068}$ $2.062^{\pm.079}$
MMM (Pinyoanuntapong et al., 2024) $0.504^{\pm.003}$ $0.696^{\pm.003}$ $0.794^{\pm.002}$ $0.080^{\pm.003}$ $2.998^{\pm.007}$ $9.411^{\pm.058}$ $1.164^{\pm.041}$
Continuous (raw data) MotionDiffuse (Zhang et al., 2024) $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
MDM (Tevet et al., 2023) $0.320^{\pm.005}$ $0.498^{\pm.004}$ $0.611^{\pm.007}$ $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $2.799^{\pm.072}$
Fg-T2M (Wang et al., 2023) $0.492^{\pm.002}$ $0.683^{\pm.003}$ $0.783^{\pm.002}$ $0.243^{\pm.019}$ $3.109^{\pm.007}$ $9.278^{\pm.072}$ $1.614^{\pm.049}$
Continuous (latent) MLD (Chen et al., 2023) $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $2.413^{\pm.079}$
MoLA (ours) $0.514^{\pm.003}$ $0.705^{\pm.002}$ $0.797^{\pm.002}$ $0.362^{\pm.009}$ $3.162^{\pm.008}$ $9.672^{\pm.079}$ $2.355^{\pm.121}$
Table 4. Comparison with state-of-the-art methods on HumanML3D.
4.3. Effectiveness of Adversarial Training
We first discuss the effect of adversarial training on Stage 1. Here, we adopt VAE and VQ-VAE described in Section 4.2 as the baselines. The case of VAE is intended to show the effects in a continuous latent space, while the case of VQ-VAE is intended to show the effects in a discrete latent space. The results for VAE combined with GAN/SAN are shown in Table 2, and those of VQ-VAE in Table 3. We utilize the reconstruction FID (rFID) as an evaluation metric. As shown in the tables, the adversarial training is effective in both VAE and VQ-VAE in the motion reconstruction task. In particular, we can improve the performance of the Stage 1 model by adopting the SAN framework instead of the conventional GAN. A better rFID is directly related to the upper bound of the overall performance of a text-to-motion model. Thus, we use VAE trained with the SAN framework for our MoLA. We demonstrate that adversarial training in Stage 1 actually improves the generation performance of the entire model in the next section.
[…]
4.4. Motion Generation
Next, we evaluate inference cost and quality of generation in comparison to existing text-to-motion methods that have publicly available implementations (Tevet et al., 2023; Chen et al., 2023; Zhang et al., 2023b; Pinyoanuntapong et al., 2024; Guo et al., 2023). Since our requirements are 1) fast generation, 2) high generation performance, and 3) training-free editing, we apply adversarial training to a continuous latent case, specifically MLD (Chen et al., 2023) (introduced in Section 3) We show that our proposed strategy improves MLD and achieves a better performance in Figure 3. Note that the difference between the MLD baseline and our MoLA is the training method, not the inference. Thus, as shown in Figure 3, inference with our model can be achieved with equivalent time to the MLD baseline, which is much faster in generation than data-space diffusion MDM (Tevet et al., 2023) that is the only existing method that can perform training-free motion editing. It is true that most methods based on discrete representation perform very well in terms of fidelity and speed, but there is no discrete method that can deal with multiple editing tasks in a training-free manner and satisfies our requirements.
In addition, we evaluate our proposed MoLA in comparison to current SOTA methods (Zhang et al., 2024; Tevet et al., 2023; Wang et al., 2023; Zhong et al., 2023; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024; Zhang et al., 2023b; Chen et al., 2023) using five metrics (R-Precision, Fréchet Inception Distance (FID), Mulit-modal distance (MMDist), Diversity, and MultiModality (MModality)) proposed by Guo et al. (Guo et al., 2022). For evaluation, we select the model that achieves the best FID, which is a metric that evaluates the overall motion quality, on the validation set and report its performance on the test set on each of the two datasets. We show the results in Tables 4 and 5. The methods are organized into three groups: i) those using VQ-based latent representations (Discrete), ii) those using data-space diffusion model (Continuous (raw data)), and iii) those using VAE-based latent representations (Continuous (latent)). The discrete approaches perform well in motion generation (e.g., (Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024)). However, those models cannot control an arbitrary set of joints in a training-free manner (Pinyoanuntapong et al., 2024) as we have discussed so far. MDM (Tevet et al., 2023), MotionDiffuse (Zhang et al., 2024) and Fg-T2M (Wang et al., 2023) adopt data-space diffusion, while MLD (Chen et al., 2023) and our MoLA utilize a diffusion model in a low-dimensional latent space. Therefore, the former are grouped in Continuous (raw data) and the latter in Continuous (latent) in the tables. As shown in Figure 3, MLD and our MoLA can be much faster in generation than MDM, which is the only existing method that can perform training-free motion editing. Moreover, MoLA achieves a better performance than MLD, especially in the R-precision, FID, and MMDist metrics on both HumanML3D and KIT-ML datasets as shown in the tables. As we will show in the next section, our MoLA can also deal with multiple motion editing tasks without additional training, which means our MoLA satisfies our three requirements best.
Category Method R-Precision $\uparrow$ FID $\downarrow$ MMDist $\downarrow$ Diversity $\rightarrow$ MModality $\uparrow$
Top-1 Top-2 Top-3
N/A Real motion data $0.424^{\pm.005}$ $0.649^{\pm.006}$ $0.779^{\pm.006}$ $0.031^{\pm.004}$ $2.788^{\pm.012}$ $11.08^{\pm.097}$
Discrete M2DM (Kong et al., 2023) $0.416^{\pm.004}$ $0.628^{\pm.004}$ $0.743^{\pm.004}$ $0.515^{\pm.029}$ $3.015^{\pm.017}$ $11.417^{\pm.97}$ $3.325^{\pm.37}$
AttT2M (Zhong et al., 2023) $0.413^{\pm.006}$ $0.632^{\pm.006}$ $0.751^{\pm.006}$ $0.870^{\pm.039}$ $3.039^{\pm.021}$ $10.96^{\pm.123}$ $2.281^{\pm.047}$
T2M-GPT (Zhang et al., 2023b) $0.416^{\pm.006}$ $0.627^{\pm.006}$ $0.745^{\pm.006}$ $0.514^{\pm.029}$ $3.007^{\pm.029}$ $10.921^{\pm.108}$ $1.570^{\pm.039}$
MoMask (Guo et al., 2023) $0.433^{\pm.007}$ $0.656^{\pm.005}$ $0.781^{\pm.005}$ $0.204^{\pm.011}$ $2.779^{\pm.022}$ $1.131^{\pm.043}$
DiverseMotion (Lou et al., 2023) $0.416^{\pm.005}$ $0.637^{\pm.008}$ $0.760^{\pm.011}$ $0.468^{\pm.098}$ $2.892^{\pm.041}$ $10.873^{\pm.101}$ $2.062^{\pm.079}$
MMM (Pinyoanuntapong et al., 2024) $0.381^{\pm.005}$ $0.590^{\pm.006}$ $0.718^{\pm.005}$ $0.429^{\pm.019}$ $3.146^{\pm.019}$ $10.633^{\pm.097}$ $1.105^{\pm.026}$
Continuous (raw data) MotionDiffuse (Zhang et al., 2024) $0.417^{\pm.004}$ $0.621^{\pm.004}$ $0.739^{\pm.004}$ $1.954^{\pm.062}$ $2.958^{\pm.005}$ $11.10^{\pm.143}$ $0.730^{\pm.013}$
MDM (Tevet et al., 2023) $0.164^{\pm.004}$ $0.291^{\pm.004}$ $0.396^{\pm.004}$ $0.497^{\pm.021}$ $9.191^{\pm.022}$ $10.847^{\pm.109}$ $1.907^{\pm.214}$
Fg-T2M  (Wang et al., 2023) $0.418^{\pm.005}$ $0.626^{\pm.004}$ $0.745^{\pm.004}$ $0.571^{\pm.047}$ $3.114^{\pm.015}$ $10.93^{\pm.083}$ $1.019^{\pm.029}$
Continuous (latent) MLD (Chen et al., 2023) $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.007}$ $0.404^{\pm.027}$ $3.204^{\pm.027}$ $10.80^{\pm.117}$ $2.192^{\pm.071}$
MoLA (ours) $0.401^{\pm.004}$ $0.612^{\pm.004}$ $0.728^{\pm.004}$ $0.350^{\pm.020}$ $3.187^{\pm.027}$ $10.88^{\pm.110}$ $2.226^{\pm.097}$
Table 5. Comparison with state-of-the-art methods on KIT-ML.
4.5. Motion Editing
Here, we demonstrate three different types of editing tasks with a single framework: editing in the time direction (in-betweening), editing in some joints of motion (upper body editing), and editing with the path specified as the trajectory (path following). We experiment with the following motion editing on models trained in HumanML3D (Guo et al., 2022).
Figure 4. Qualitative results of motion in-betweening task. Blue and red points are used as inputs to constrain the start and end positions, respectively. These are the results generated by our framework from the same text prompt (“A person walks while raising his left hand.”) for two different start-end positions.
In-betweening: Motion in-betweening is an important editing task that interpolates or fills the gaps between keyframes or major motion joints to create smooth 3D motion animation. In our framework, by applying the update rule of MPGD (He et al., 2024) on latent space (Section 3.4), it is easy to coordinate and generate motion between past and future contexts without additional training. We only need to set the start-end positions or the motions of a few frames as the control signal $\bm{y}$ in (12). The motion in-betweening results with our model when different start-end controls are given as $\bm{y}$ in the same text condition are shown in Figure 4.
Figure 5. Qualitative results of upper body editing task. Each colored point is used inputs to control the lower body position as a constraint. These are the results generated by our framework from the same text prompt (“A person puts his right hand to the air.”) for two different lower body positions.
Upper body editing: Upper body editing combines generated upper body parts with given lower body parts. Generating certain joints to fit given body parts can be seen as the task of outpainting in the spatial dimension of motion. In our framework, similar to the motion in-betweening, the loss and its gradient are calculated in the data-space as (11), and sampling is performed taking into account its gradients. The control signal $\bm{y}$ in (12) is set as lower body positions that is not subject to editing. Note that although we are dealing with upper body editing in these experiments, it is in principle possible to specify a different joint subset. Therefore, without the need for a separate architecture or additional training, our framework allows editing conditioned on given body parts of motion. The upper body editing results with our model when different lower body controls are given as $\bm{y}$ in the same text condition are shown in Figure 5.
Figure 6. Qualitative results of path-following task. The points are the input to control a motion trajectory. These are the results generated by our framework from the same text prompt (“A person walks with his hands up.”) for two different given trajectories.
Path-following: Path-following motion generation is a task of giving a trajectory (often the position of the pelvis) and generating the motion that matches with the given route. Controlling the trajectory of generated motion can lead to more motion variation, avoidance of obstacles, and the creation of motion that meets physical constraints. Our framework can edit a motion to follow a specific path by minimizing the loss between a joint of the generated motion and a desired trajectory of the joint, similar to motion in-betweening. In this experimental case, the desired pelvis trajectory is set as the control signal $\bm{y}$ in (12). The path-following generation results with our model when different path controls are given as $\bm{y}$ in the same text condition are shown in Figure 6.
5. Conclusion
In this paper, we proposed MoLA, a text-to-motion model that achieves fast, high-quality generation with multiple control tasks in a single framework. To construct our model, we employed a latent diffusion-based method, and introduced adversarial training. In addition, we applied a training-free guided generation framework based on MPGD. The results of quantitative evaluation demonstrate the effectiveness of adversarial training in the text-to-motion generation. In particular, we found that replacing GAN with SAN boosts the performance of our model. We also showed that our model can perform motion in-betweening, upper body editing, and path following within a single framework.
Appendix A Notation
In Table 6, we list the several notations and symbols that we have used in this paper.
Notation/Symbol Description
$\mathbb{N}$ The sets of natural numbers
$\mathbb{R}$ The sets of real numbers
$\mathbb{R}_{+}$ The sets of positive real numbers
$N\in\mathbb{N}$ A motion feature dimension
$L\in\mathbb{N}$ A motion length
$d_{z}\in\mathbb{N}$ A latent space dimension
$d_{c}\in\mathbb{N}$ A conditional embedding dimension
$t\in\mathbb{N}$ A time step for diffusion model
$\bm{\eta}$ A trainable parameter for the encoder
$\bm{\psi}$ A trainable parameter for the decoder
$\bm{\phi}=\{\bm{\varphi},\bm{\omega}\}$ A trainable parameter for the discriminator
$\bm{\theta}$ A trainable parameter for the estimator in latent diffusion
$\bm{x}\in\mathbb{R}^{N\times L}$ A motion input
$\bm{z}\in\mathbb{R}^{d_{z}}$ A latent variable
$\tau(\bm{c})\in\mathbb{R}^{d_{c}}$ A output of text encoder corresponding to a prompt $\bm{c}$
$q_{\bm{\eta}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{z}}\times\mathbb{R}_{+}^{d_{z}}$ A encoder parameterized by $\bm{\eta}$ in Stage 1
$g_{\bm{\psi}}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{N\times L}$ A decoder parameterized by $\bm{\psi}$ in Stage 1
$f_{\bm{\phi}}:\mathbb{R}^{N\times L}\to\mathbb{R}$ A discriminator parameterized by $\bm{\phi}$ in Stage 1
$\epsilon_{\bm{\theta}}:\mathbb{R}^{d_{z}}\times\mathbb{N}\times\mathbb{R}^{d_{c}}\to\mathbb{R}^{d_{z}}$ A conditional estimator parameterized by $\bm{\theta}$ in Stage 2
$\lambda_{\text{reg}}\in\mathbb{R}_{+}$ A weight of VAE regularization term in Stage 1
$\lambda_{\text{adv}}\in\mathbb{R}_{+}$ A weight of adversarial loss in Stage 1
$\alpha_{t}\in(0,1)$ A noise scheduling parameter in Stage 2
$\rho_{t}\in\mathbb{R}_{+}$ A time-dependent step size in guided generation
Table 6. Notation table
Appendix B Editing procedure on guided generation
In Algorithm 1, we show the specific procedure for guided generation described in Section 3.4. Note that we apply a time-travel technique to the update rule in Section 3.4 as in (Lugmayr et al., 2022; Wang et al., 2022; Yu et al., 2023; He et al., 2024) to achieve better editing results. This technique adds noise after each gradient descent step to implicitly perform a multi-step optimization of minimizing the distance measuring function $L(\cdot,\bm{y})$ and lead to an improvement in the editing quality.
Algorithm 1 Guided Generation for Motion Editing
0:  motion control signal $\bm{y}$, output of text encoder $\tau(\bm{c})$, estimator in latent diffusion $\epsilon_{\theta}(\cdot,t,\tau(\bm{c}))$, distance measuring function $L(\cdot;\bm{y})$, pre-defined parameter $\bar{\alpha}_{t}$, time-dependent step-size $\rho_{t}$, and the repeat times of time-travel of each step $\{r_{1},\cdots,r_{T}\}$.
1:  $\bm{z}_{T}\sim\mathcal{N}(0,\bm{I})$
2:  for $t=T,\cdots,1$ do
3:     for $i=r_{t},\cdots,1$ do
4:        $\epsilon_{t}\sim\mathcal{N}(0,\bm{I})$
5:        $\bm{z}_{0|t}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c})))$
6:        $\bm{z}_{0|t}=\bm{z}_{0|t}-\rho_{t}\nabla_{\bm{z}_{0|t}}L(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y})$
7:        $\bm{z}_{t-1}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0|t}+\sqrt{1-\bar{\alpha}_{t}-\sigma^{2}_{t}}\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))+\sigma_{t}\epsilon_{t}$
8:        if $i>1$ then
9:           $\epsilon_{t}^{\prime}\sim\mathcal{N}(0,\bm{I})$
10:           $\bm{z}_{t}=\sqrt{\alpha_{t}}\bm{z}_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t}^{\prime}$
11:        end if
12:     end for
13:  end for
14:  return  $g_{\bm{\psi}}(\bm{z}_{0})\qquad\qquad\triangleright$ edited motion generation sample
Methods FID $\downarrow$ R-Precision Top-1 $\uparrow$ MMDist $\downarrow$
T2M-GPT (Zhang et al., 2023b) $0.141^{\pm.005}$ $0.492^{\pm.003}$ $3.121^{\pm.009}$
T2M-GPT + GAN (ours) $0.136^{\pm.010}$ $0.485^{\pm.004}$ $3.116^{\pm.009}$
T2M-GPT + SAN (ours) $0.133^{\pm.006}$ $0.494^{\pm.003}$ $3.111^{\pm.013}$
MLD (Chen et al., 2023) $0.473^{\pm.013}$ $0.481^{\pm.003}$ $3.196^{\pm.010}$
MLD + GAN (ours) $0.371^{\pm.014}$ $0.482^{\pm.002}$ $3.173^{\pm.011}$
MLD + SAN (MoLA; ours) $0.362^{\pm.009}$ $0.514^{\pm.003}$ $3.162^{\pm.008}$
Table 7. Effectiveness of adversarial training for motion generation performance on HumanML3D
Methods FID $\downarrow$ R-Precision Top-1 $\uparrow$ MMDist $\downarrow$
T2M-GPT (Zhang et al., 2023b) $0.514^{\pm.029}$ $0.416^{\pm.006}$ $3.007^{\pm.029}$
T2M-GPT + GAN (ours) $0.484^{\pm.036}$ $0.408^{\pm.008}$ $3.027^{\pm.022}$
T2M-GPT + SAN (ours) $0.439^{\pm.033}$ $0.423^{\pm.006}$ $2.947^{\pm.030}$
MLD (Chen et al., 2023) $0.404^{\pm.027}$ $0.390^{\pm.008}$ $3.204^{\pm.027}$
MLD + GAN (ours) $0.382^{\pm.028}$ $0.385^{\pm.005}$ $3.305^{\pm.035}$
MLD + SAN (MoLA; ours) $0.350^{\pm.020}$ $0.401^{\pm.004}$ $3.187^{\pm.027}$
Table 8. Effectiveness of adversarial training for motion generation performance on KIT-ML
Appendix C Evaluation metrics details
We provide more details of evaluation metrics in Section 4.4. We use five metrics to quantitatively evaluate text-to-motion models. These metrics are calculated based on motion and text features extracted with pre-trained networks. More specifically, we utilize the motion encoder and text encoder provided in (Guo et al., 2022). We denote ground-truth motion features, generated motion features, and text features as $f_{\text{gt}}=\mathcal{E}_{\text{M}}(x_{\text{gt}})\in\mathbb{R}^{512}$, $f_{\text{pred}}=\mathcal{E}_{\text{M}}(x_{\text{pred}})\in\mathbb{R}^{512}$, and $f_{\text{text}}=\mathcal{E}_{\text{T}}(c)\in\mathbb{R}^{512}$, where $\mathcal{E}_{\text{M}}(\cdot)$ and $\mathcal{E}_{\text{T}}(\cdot)$ represent the motion encoder and the text encoder, respectively. We explain the five metrics below.
R-Precision: R-Precision evaluates semantic alignment between input text and generated motion in a sample-wise manner. Given one motion sequence and 32 text descriptions (one ground truth and thirty-one randomly selected mismatched descriptions), we rank the Euclidean distances between the motion and text embeddings. Then, we compute the average accuracy at top-1, top-2, and top-3 places.
Fréchet Inception Distance (FID): FID evaluates the overall motion quality by measuring the distributional difference between the motion features of the generated motions and those of real motions. We obtain FID by
(13) $\text{FID}:=\|\mu_{\text{gt}}-\mu_{\text{pred}}\|^{2}_{2}-\mathrm{tr}(\Sigma_{\text{gt}}+\Sigma_{\text{pred}}-(2\Sigma_{\text{gt}}\Sigma_{\text{pred}})^{\frac{1}{2}}),$
where $\mu_{\text{gt}}$ and $\mu_{\text{pred}}$ are the mean of $f_{\text{gt}}$ and $f_{\text{pred}}$, respectively, $\Sigma_{\text{gt}}$ and $\Sigma_{\text{pred}}$ are their corresponding covariance matrices, and $\mathrm{tr}$ denotes the trace of a matrix.
Multi-modal distance (MMDist): MMDist gauges sample-wise semantic alignment between input text and generated motion as well. MMDist is computed as the average Euclidean distance between the motion feature of each generated motion and the text feature of its corresponding description in the test set. Precisely, given $N$ randomly generated samples, it computes the average Euclidean distance between each text feature and its corresponding generated motion feature:
(14) $\text{MMDist}:=\frac{1}{N}\sum_{i=1}^{N}\|f_{\text{pred},i}-f_{\text{text},i}\|_{2}^{2},$
where $f_{\text{pred},i}$ and $f_{\text{text},i}$ are the features of the $i$-th text-motion pair.
Diversity: Diversity measures the variance of the generated motions across the test set. We randomly sample $2S_{d}$ motions and extract motion features $\{f_{\text{pred},1},...,f_{\text{pred},2S_{d}}\}$ from the motions. Then, those motion features are grouped into the same size of two subsets, $\{v_{1},...,v_{S_{d}}\}$ and $\{v^{\prime}_{1},...,v^{\prime}_{S_{d}}\}$. The Diversity for this set of motions is defined as
(15) $\text{Diversity}:=\frac{1}{S_{d}}\sum_{i=1}^{S_{d}}\|v_{i}-v^{\prime}_{i}\|_{2}^{2}.$
$S_{d}=300$ is used in our experiments.
Multi-modality (MModality): MModality measures how much the generated motions diversify within each text description. Given a set of motions with $K$ text descriptions, we randomly sample $2S_{l}$ motions corresponding to the $k$-th text description and extract motion features $\{f_{\text{pred},k,1},...,f_{\text{pred},k,2S_{l}}\}$. Then, those motion features are grouped into the same size of two subsets, $\{v_{k,1},...,v_{k,S_{l}}\}$ and $\{v^{\prime}_{k,1},...,v^{\prime}_{k,S_{l}}\}$. The MModality for this motion set is formalized as
(16) $\text{MModality}=\frac{1}{K\cdot S_{l}}\sum_{k=1}^{K}\sum_{i=1}^{S_{l}}\|v_{k,i}-v^{\prime}_{k,i}\|_{2}^{2}.$
$S_{l}=10$ is used in our experiments.
Appendix D Additional experiments for motion generation
Here, we discuss the effect of adversarial training on Stage 2 to supplement the experiments in Sections 4.3 and 4.4. To demonstrate the broad effectiveness of adversarial training in text-to-motion generation, we apply adversarial training not only to a continuous latent case as in MLD (Chen et al., 2023) introduced in Section 3 but also to a discrete latent case as in T2M-GPT (Zhang et al., 2023b) introduced in Sections 4.2 and 4.3. The results of two types of methods, (top:) a discrete latent case and (bottom:) a continuous latent case, are shown in Tables 7 and 8. As shown in the tables, our proposed strategy improves existing methods in FID, which is a metric that evaluates the overall motion quality, on both continuous latent and discrete latent cases on both HumanML3D (Guo et al., 2022) and KIT-ML (Plappert et al., 2016) datasets. Moreover, we can see replacing GAN with SAN boosts performance more on FID, R-precision, and MMDist metrics on both datasets. Note that the only difference between baseline models and our models is training method, not inference, as mentioned in Section 4.4. Therefore, inference with our model can be achieved with equivalent time to MLD and T2M-GPT baselines.